Suppose d() = 8% . Find equivalent rates d , d), i , and i).

answer 1 and 2
Suppose d() = 8% . Find equivalent rates d , d), i , and ;). Suppose d(4) = 3.2% . Find 8 .

answer 1 and 2
Answer #1

0 () 846 Lida (1-11 - day) a) - 1-C1-0.0%)? do 0.0776318.4 -> d= 7.76% des) 361-ci- 0.08.)))= 0.079 73213 des) 0.079732.13) -☺ drø– 3.2%, to find s. 8 = ln (1+i) = ln (C1 - do)-9) = - 4 ln (1- - = 0.032128687 - A in (1 - 0032 (8 = 0.032128687 that 8
Answer #2

© Given de 8% do 1-C1-2008) 4 = 0.07763184 =7.7677 (d) 3 = 3 (1-11-08) 0.07973 (7.977 is(-0.05) 4., . = 0.084165785 = 5.40167G Given 3 3.2 % da ln (1+i) = ln (1-2) =-4 ln (1-day) =-4 ln (1-0-03-2) = 0.0321287 [5= 3.21%